OpenStudy (lilai3):

Find the value of r so the line that passes through each pair of points has the give slope. (4,r), (r,2), m=-5/3 (fractional form) the formula is y2-y1/ x2-x1 (fractional form) THANKS Y'ALL!!! I'LL REWARD MEDAL/ FAN YOU!!!! ~~

OpenStudy (anonymous):

now just set up your equation using the formula (substituting in your given points) and set it equal to the desired slope

OpenStudy (lilai3):

yes, but i redid it two times. and however, both times i got a totally different answer, so yeah. that's what i need help with. LOL. the solving part yeh

OpenStudy (anonymous):

\[\frac{2-r }{ r-4 }=\frac{ -5 }{ 3 }\] cross multiply and get 3(2-r)=-5(r-4) distribute 6-3r=-5r+20 then combine like terms ect to solve for r

OpenStudy (anonymous):

the important thing is to make sure you keep the order of your points straight when you plug them into the equation. A common mistake is to do \[\frac{ y2-y1 }{x1-x2 }\] which is wrong it doesn't matter if you do y1 or y2 first as long as the denominator mirrors the order of the numerator

OpenStudy (lilai3):

is the answer 0?

OpenStudy (anonymous):

if you substitute 0 into the equation does it make sense? 6-3(0)=-5(0)+20

OpenStudy (lilai3):

no it's not equal

OpenStudy (anonymous):

okay so if we combine like terms in the equation 6-3r=-5r+20 we get that 2r=14 do you know how to do that?

OpenStudy (lilai3):

so r equals 7?

OpenStudy (anonymous):

yep =) but more importantly, do you know how to do the problem yourself? which part messed you up?

OpenStudy (lilai3):

the -19/17 part

OpenStudy (anonymous):

hmm where did the -19/17 come in at? lol

OpenStudy (lilai3):

oh...it wasn't there? can you show me how you got your answer of 7?

OpenStudy (lilai3):

@lmarl

OpenStudy (anonymous):

start with your slope equation m=(y2-y1)/(x2-x1) then sub in your points: m=(2-r)/(r-4) then because we want the slope to be -5/3 we set m in the equation to that value: -5/3=(2-r)/(r-4) now we cross multiply to get rid of the fractions: 3(2-r)=-5(r-4) then distribute: 6-3r=-5r+20 now move all the variables to one side and all the constants to the other: -3r+5r=20-6 now combine like terms: 2r=14 and divide r=7 =)